### v1

@@ Abstract
We establish a unified synchronization framework for the all-to-all hybrid Kuramoto model that couples first- and second-order oscillators within a single dynamical system. Although the Kuramoto model has become one of the most widely used paradigms for describing synchronization phenomena-appearing in more than $100,000$ scientific studies-the fundamental relationships among distinct synchronization states remain unresolved. In this work, we rigorously prove that full phase-locking, phase-locking, frequency synchronization, and order-parameter synchronization are equivalent for arbitrary hybrid ensembles. The proof combines dissipative energy methods, LaSalle-type compactness arguments, the Poincaré-Bendixson theorem, and Thieme’s asymptotically autonomous theory to demonstrate that synchronization equivalence is topological, determined solely by the finite equilibrium structure of the all-to-all network. This result provides a mathematically complete characterization of synchronization in finite oscillator systems and clarifies its geometric invariance across first-, second-, and hybrid-order models.
@@ 2.1 Definitions of Synchronization State
Definition 2.1 (PS).
A solution $\theta(t)$ is called phase synchronization state if for all $j,k\in\{1,\ldots,N\}$,
@@
A solution $\theta(t)$ is called full phase-locked state if for all $j,k\in\{1,\ldots,N\}$,
@@
A solution $\theta(t)$ is called phase-locked state if for all $j,k\in\{1,\ldots,N\}$, there exists a positive constant $L$ such that for all $t>0$,
@@
A solution $\theta(t)$ is called frequency synchronization state if
@@
$\lim_{t\rightarrow\infty}Z(t)=Z^{*}\quad\text{and}\quad\left|Z^{*}\right|\geq\frac{\left|\omega_{j}\right|}{\lambda}~~\mbox{for}~\mbox{all}~j\in\{1,\ldots,N\}.$ (2.2)
We pause to remark that a necessary condition for the existence of a phase synchronization state is that all natural frequencies are identical. In other words, one must assume that
@@ 2.2 Main Theorems
$\theta_{m}=\min_{1\leq j\leq N}\theta_{j},\qquad\theta_{M}=\max_{1\leq j\leq N}\theta_{j},$ (2.4)
$\omega_{M}=\max_{1\leq j\leq N}\omega_{j},\qquad\omega_{m}=\min_{1\leq j\leq N}\omega_{j}.$
@@
Consequently, any of the above equivalent properties may be used as the definition of a synchronization state.
@@
$\sin(\Psi-\psi_{j})=-\frac{\omega_{j}}{\lambda r}:=s_{j},\quad\text{with}\quad r\in\left[\frac{\max\{\omega_{M},|\omega_{m}|\}}{\lambda},1\right].$
@@
$rN=\sum_{j=1}^{N}\cos(\Psi-\psi_{j})=\sum_{j=1}^{N}\sigma_{j}\sqrt{1-\left(\frac{\omega_{j}}{\lambda r}\right)^{2}},\,\text{with}\quad r\in\left[\frac{\max\{\omega_{M},|\omega_{m}|\}}{\lambda},1\right].$
Fix a sign vector $\sigma=(\sigma_{1},\dots,\sigma_{N})\in\{\pm 1\}^{N}$ and set $x=\lambda r\in[\max\{\omega_{M},|\omega_{m}|\},\,\lambda]$. One may define
@@
Since $H_{\sigma}$ is a nontrivial real-analytic function on $(\omega_{M},\infty)$, its zeros are isolated. Hence $H_{\sigma}(x)=0$ has at most finitely many solutions within the compact interval $\max\{\omega_{M},|\omega_{m}|\},\lambda]$. Because there are at most $2^{N}$ choices of $\sigma$, it follows that the total number of admissible values of $r$ is finite. For each fixed pair $(r,\sigma)$, the angle
@@
Let $I\subset\mathbb{R}^{+}$ be an open set and let $f\in C^{2}(I)$ satisfy $\|f\|_{L^{\infty}(I)}<\infty$ and $\|\ddot{f}\|_{L^{\infty}(I)}<\infty$. Then
@@ 3.2 Proof of $\mathrm{(SS)}\Leftrightarrow\mathrm{(OPSS)}$
$|\omega_{k}|>R(t)|\lambda|,\quad\text{for }t\text{ sufficiently large}.$
@@
We pause to remark that, since the system is $2\pi$-periodic, the vector field is invariant under the transformation $\theta\mapsto\theta+2\pi$. Therefore, we first consider the dynamics projected onto the torus $\mathbb{T}^{N}$ and subsequently lift the trajectories back to the original space $\mathbb{R}^{N}$.
@@
Suppose, for contradiction, that a periodic orbit exists. By the Bendixson–Dulac theorem, any contractible closed orbit in $\mathbb{R}\times\mathbb{T}$ cannot occur, since the divergence satisfies
@@
whose sign is strictly negative and never vanishes. Hence, the only remaining possibility for a periodic orbit (with period $T$) is a non-contractible one winding around the cylinder $\mathbb{R}\times\mathbb{T}$.
Let us lift this periodic solution back to $\mathbb{R}\times\mathbb{R}$ and consider the Lyapunov-type function
@@
$\frac{d}{dt}L(v_{j}(t),\theta_{j}(t))=-d_{j}v_{j}^{2}\leq 0.$ (3.15)
We begin with a preliminary observation. Suppose the orientation around the cylinder is taken to be positive; that is, when lifted to $\mathbb{R}^{2}$, $\theta_{j}$ increases toward $\infty$. After one period $T$, we have $v_{j}(T)=v_{j}(0)$ and $\theta_{j}(T)=\theta_{j}(0)+2\pi$. Hence, integrating (3.15) over $[0,T]$ and using (3.14), we obtain
$2\pi\omega_{j}-d_{j}\int_{0}^{T}v_{j}^{2}(t)\,dt=0,$
which implies $\omega_{j}>0$ is a necessary condition. Conversely, if $\theta_{j}(t)$ tends to -$\infty$, then after one period $T$, we have $v_{j}(T)=v_{j}(0)$ and $\theta_{j}(T)=\theta_{j}(0)-2\pi$. Similarly, we have
$2\pi\omega_{j}+d_{j}\int_{0}^{T}v_{j}^{2}(t)\,dt=0,$
hence $\omega_{j}<0$.
Consequently, if a periodic orbit winding around the cylinder exists, it suffices to consider the case $\omega_{j}>0$ (and $\theta_{j}\to+\infty$), or equivalently $\omega_{j}<0$ with $\theta_{j}\to$-$\infty$ when lifted to $\mathbb{R}^{2}$. In what follows, we focus on the former case.
@@
Fix $\theta_{j}(0)=\theta_{0}$ and regard the initial speed $v_{0}>0$ as a free parameter. Define a Poincaré-like map $P$ so that $P(v_{0})\in\Sigma^{+}$; namely, starting from the initial condition $(v_{j}(0),\theta_{j}(0))=(v_{0},\theta_{0})$, the value $P(v_{0})$ represents the velocity component of the solution upon its first return to the cross section $\Sigma^{+}$. If $P$ admits a fixed point, then a periodic orbit exists. Observe that for $v_{0}=0$ we obtain the stationary solution $(v_{j},\theta_{j})\equiv(0,\theta_{0})$. Let $\tau=\tau(v_{0})$ denote the (first) return time required for the trajectory with initial speed $v_{0}$ to intersect $\Sigma^{+}$.
By (3.14) and (3.15), we have
@@
Hence, no positive fixed point can exist, and consequently no periodic orbit $(v_{j},\theta_{j})\in\mathbb{R}\times\mathbb{T}$ arises. So far, by applying Theorem 3.3, we can only guarantee that for any $\gamma$-orbit (which is automatically pre-compact) of (3.12), the trajectory $\gamma(t,s;v_{j},\theta_{j})$ either converges to an equilibrium point of (3.13), or the $\omega$-$\gamma$-limit sets of $(s,v_{j},\theta_{j})$ contain finitely many equilibria of (3.13) that are chained together in a cyclic way. Therefore, our next goal is to analyze the existence of homoclinic and heteroclinic orbits in (3.13).
We exclude the possibility of non-contractible homoclinic or heteroclinic orbits. If such orbits existed, by an argument analogous to the previous one, when lifted to $\mathbb{R}\times\mathbb{R}$, they would become a homoclinic/heteroclinic ladder that advances monotonically in the $\theta_{j}$-direction:
@@
$(v_{j}(t_{k}),\theta_{j}(t_{k}))=(v_{j}(t_{k}),\theta_{0}+2k\pi)\quad\text{and}\quad v_{j}(t_{k})\to 0\quad\text{as}\quad t_{k}\to\infty.$
However, by applying the previous argument $P(v_{j}(t_{k}))>v_{j}(t_{k})$ (cf. (3.20)), we conclude that such a situation cannot occur. The only remaining possibility is the existence of contractible homoclinic or heteroclinic orbits. This case is “mild,” because when we lift the system back to $\mathbb{R}\times\mathbb{R}$, the cyclic chain formed by these contractible homoclinic or heteroclinic orbits remains bounded. If the $\omega$-$\gamma$-limit sets of $(s,v_{j},\theta_{j})$ contain finitely many equilibria of (3.13) connected by such bounded homoclinic or heteroclinic orbits, then we know that $\gamma$ remains bounded (even after lifting to $\mathbb{R}\times\mathbb{R}$). Therefore, we conclude that such trajectories must eventually become (SS). We complete the proof. ∎
@@ Acknowledgements
Data will be made available on reasonable request.

### v2

@@ Abstract
We establish a unified synchronization framework for the all-to-all hybrid Kuramoto model that couples first- and second-order oscillators within a single dynamical system. Although the Kuramoto model has become one of the most widely used paradigms for describing synchronization phenomena-appearing in more than 100,000 scientific studies-the fundamental relationships among distinct synchronization states remain unresolved. In this work, we rigorously prove that full phase-locking, phase-locking, frequency synchronization, and order-parameter synchronization are equivalent for arbitrary hybrid ensembles. The proof combines dissipative energy methods, LaSalle-type compactness arguments, the Poincaré-Bendixson theorem, and Thieme’s asymptotically autonomous theory to demonstrate that synchronization equivalence is topological, determined solely by the finite equilibrium structure of the all-to-all network. This result provides a complete mathematical characterization of synchronization in finite oscillator systems and clarifies its geometric invariance across first-, second-, and hybrid-order models.
@@ 2.1 Definitions of Synchronization State
Definition 2.1 (PSS).
A solution $\theta(t)$ is called a phase synchronization state if for all $j,k\in\{1,\ldots,N\}$,
@@
A solution $\theta(t)$ is called a full phase-locked state if for all $j,k\in\{1,\ldots,N\}$,
@@
A solution $\theta(t)$ is called a phase-locked state if for all $j,k\in\{1,\ldots,N\}$, there exists a positive constant $L$ such that for all $t>0$,
@@
A solution $\theta(t)$ is called a frequency synchronization state if
@@
$\lim_{t\rightarrow\infty}Z(t)=Z^{*}\quad\text{and}\quad\left|Z^{*}\right|\geq\frac{\omega_{M}}{\lambda},$ (2.2)
where $\omega_{M}:=\max\limits_{j\in\{1,\ldots,N\}}|\omega_{j}|$. We pause to remark that a necessary condition for the existence of a phase synchronization state (PSS) is that all natural frequencies are identical. In other words, one must assume that
@@ 2.2 Main Theorems
$\theta_{m}=\min_{1\leq j\leq N}\theta_{j},\qquad\theta_{M}=\max_{1\leq j\leq N}\theta_{j},\qquad\omega_{M}=\max_{1\leq j\leq N}\left|\omega_{j}\right|.$ (2.4)
@@
Consequently, any of the above equivalent properties may be used as the definition of a synchronization state. Moreover, in the special case $\omega_{j}=0$, the previous synchronization notions are equivalent to $\mathrm{(PSS)}$ as well.
@@
$\sin(\Psi-\psi_{j})=-\frac{\omega_{j}}{\lambda r}:=s_{j},\quad\text{with}\quad r\in\left[\frac{\omega_{M}}{\lambda},1\right].$
@@
$rN=\sum_{j=1}^{N}\cos(\Psi-\psi_{j})=\sum_{j=1}^{N}\sigma_{j}\sqrt{1-\left(\frac{\omega_{j}}{\lambda r}\right)^{2}},\,\text{with}\quad r\in\left[\frac{\omega_{M}}{\lambda},1\right].$
Fix a sign vector $\sigma=(\sigma_{1},\dots,\sigma_{N})\in\{\pm 1\}^{N}$ and set $x=\lambda r\in[\omega_{M},\lambda]$. One may define
@@
Since $H_{\sigma}$ is a nontrivial real-analytic function on $(\omega_{M},\infty)$, its zeros are isolated. Hence $H_{\sigma}(x)=0$ has at most finitely many solutions within the compact interval $\omega_{M},\lambda]$. Because there are at most $2^{N}$ choices of $\sigma$, it follows that the total number of admissible values of $r$ is finite. For each fixed pair $(r,\sigma)$, the angle
@@
Let $I\subset\mathbb{R}^{+}$ be an open interval and $f\in C^{2}(I)$ satisfy $\|f\|_{L^{\infty}(I)}<\infty$ and $\|\ddot{f}\|_{L^{\infty}(I)}<\infty$. Then
@@ 3.2 Proof of $\mathrm{(SS)}\Leftrightarrow\mathrm{(OPSS)}$
$|\omega_{k}|>\left|Z(t)\right||\lambda|=R(t)|\lambda|,\quad\text{for }t\text{ sufficiently large}.$
@@
We note that since the system is $2\pi$-periodic, the vector field is invariant under the transformation $\theta\mapsto\theta+2\pi$. Therefore, we first consider the dynamics projected onto the torus $\mathbb{T}^{N}$ and subsequently lift the trajectories back to the original space $\mathbb{R}^{N}$.
@@
First, by the Bendixson–Dulac theorem, any contractible closed orbit in $\mathbb{R}\times\mathbb{T}$ cannot occur, since the divergence satisfies
@@
whose sign is strictly negative and never vanishes. Therefore, both a contractible periodic orbit and a contractible homoclinic orbit are impossible. Next, if there exist two heteroclinic orbits connecting the two equilibria, they can be further classified into two topologically distinct cases: a pair of contractible heteroclinic orbits, which together bound a simply connected region on $\mathbb{R}\times\mathbb{T}$, or a pair of non-contractible heteroclinic orbits, which wind once around the cylindrical direction of $\mathbb{R}\times\mathbb{T}$ and therefore cannot be continuously deformed into each other within the phase space. Let us get rid of the first case. By lifting this pair of contractible heteroclinic orbits back to $\mathbb{R}\times\mathbb{R}$ and consider the Lyapunov-type function
@@
$\frac{d}{dt}L(v_{j}(t),\theta_{j}(t))=-d_{j}v_{j}^{2}<0\quad\text{when}\quad v_{j}\neq 0.$ (3.15)
Therefore, if there exist two heteroclinic orbits connecting the two equilibria $e_{1}$ and $e_{2}$, the Lyapunov function $L$ must satisfy $L(e_{1}^{+})>L(e_{2}^{+})$ along one orbit and $L(e_{1}^{+})<L(e_{2}^{+})$ along the other, which is impossible.
Next, we exclude the more delicate cases of a non-contractible periodic orbit, a non-contractible homoclinic orbit, and a pair of non-contractible heteroclinic orbits. Lifting the dynamics from $\mathbb{R}\times\mathbb{T}$ to $\mathbb{R}\times\mathbb{R}$, all such trajectories become unbounded in the lifted space, regardless of whether they wind with positive or negative orientation on the cylinder. Without loss of generality, we may assume that the angular component satisfies $\limsup\limits_{t\rightarrow\infty}\theta_{j}(t)\to+\infty$. The other case can be argued by a similar method.
@@
Fix $\theta_{j}(0)=\theta_{0}$ and regard the initial speed $v_{0}>0$ as a free parameter. Define a Poincaré-like map $P$ so that $P(v_{0})\in\Sigma^{+}$; namely, starting from the initial condition $(v_{j}(0),\theta_{j}(0))=(v_{0},\theta_{0})$, the value $P(v_{0})$ represents the velocity component of the solution upon its first return to the cross section $\Sigma^{+}$. This is indeed inevitable. Since the lifted trajectory is unbounded with $\theta_{j}(t)\to+\infty$, let
$\tau(v_{0}):=\inf\{\,t>0:\ \theta_{j}(t)=\theta_{0}+2\pi\,\}$
be the first hitting time of $\theta_{j}=\theta_{0}+2\pi$. By continuity of $\theta_{j}(t)$ and monotonic escape in the angular direction, $\tau(v_{0})<\infty$. At $t=\tau(v_{0})$ we cannot have $v_{j}(\tau(v_{0}))=0$, because then, by uniqueness of solutions, the trajectory would stick to the equilibrium and be trivial, contradicting unboundedness. Nor can $v_{j}(\tau(v_{0}))<0$, since this would violate the choice of positive orientation toward $\theta_{j}\to+\infty$ at the first crossing. Therefore the Poincaré return velocity satisfies $P(v_{0}):=v_{j}(\tau(v_{0}))>0$.
Lemma 3.5.
Let $(v(t;v_{0}),\theta(t;v_{0}))$ solve (3.13) with $(v(0;v_{0}),\theta(0;v_{0}))=(v_{0},\theta_{0})$, and let $\tau(v_{0})>0$ be the first time $\theta(t;v_{0})=\theta_{0}+2\pi$, assuming $v(\tau(v_{0});v_{0})>0$. Define $P(v_{0}):=v(\tau(v_{0});v_{0})$. Then $\tau$ and $P$ are of class $C^{1}$ in a neighborhood of $v_{0}$.
Proof.
Let $G(t,v_{0})=\theta(t;v_{0})-(\theta_{0}+2\pi)$. Since $G(\tau(v_{0}),v_{0})=0$ and $\partial_{t}G=v(\tau(v_{0});v_{0})\neq 0$, the implicit function theorem gives $\tau\in C^{1}$ and the first identity. The second follows by the chain rule applied to $P(v_{0})=v(\tau(v_{0});v_{0})$. ∎
We now exclude the possibility of periodic orbits. We notice that if $P$ admits a fixed point, then a periodic orbit exists. By (3.14) and (3.15), we have
@@
Hence, no positive fixed point can exist, and consequently no periodic orbit $(v_{j},\theta_{j})\in\mathbb{R}\times\mathbb{T}$ arises.
We exclude the cases a non-contractible homoclinic orbit, and a pair of non-contractible heteroclinic orbits. If such orbits existed, by an argument analogous to the previous one, when lifted to $\mathbb{R}\times\mathbb{R}$, we obtain homoclinic/heteroclinic ladder of equilibria that advances monotonically in the $\theta_{j}$-direction:
@@
$(v_{j}(t_{k}),\theta_{j}(t_{k}))=(v_{j}(t_{k}),\theta_{0}+2k\pi).$
Because the trajectory repeatedly approaches the equilibria (homoclinic or heteroclinic ladder of equilibria), there exists a subsequence $\{t_{k_{\ell}}\}$ for which $v_{j}(t_{k_{\ell}})\to 0$ as $\ell\to\infty$. However, by applying the previous argument $P(v_{j}(t_{k_{l}}))>v_{j}(t_{k_{l}})$ (cf. (3.20)), we conclude that such a situation cannot occur.
So far, by applying Theorem 3.3, we can only guarantee that for any $\gamma$-orbit (which is automatically pre-compact) of (3.12), the trajectory $\gamma(t,s;v_{j},\theta_{j})$ either converges to an equilibrium point of (3.13), or the $\omega$-$\gamma$-limit sets of $(s,v_{j},\theta_{j})$ contain a single heteroclinic orbit connecting the two equilibria of (3.13). If only the first case occurs, the proof is complete. Also, the latter case is “mild,” because when we lift the system back to $\mathbb{R}\times\mathbb{R}$, the cyclic chain formed by these contractible homoclinic or heteroclinic orbits remains bounded. If the $\omega$-$\gamma$-limit sets of $(s,v_{j},\theta_{j})$ contain finitely many equilibria of (3.13) connected by such bounded homoclinic or heteroclinic orbits, then we know that $\gamma$ remains bounded (even after lifting to $\mathbb{R}\times\mathbb{R}$). Therefore, we conclude that such trajectories must eventually become (SS). We complete the proof. ∎
@@ Acknowledgements
This study does not involve any data.